OpenStudy (anonymous):

Please help ... Evaluate : Sqr Rt Of 81 + 3^2 - 6 - Sqr Rt Of 16 .....

OpenStudy (lgbasallote):

what is \(\sqrt{81}\)?

OpenStudy (lgbasallote):

what do you think is the value of that?

OpenStudy (anonymous):

its 5 xD

OpenStudy (anonymous):

9-4 = 5 ow im wrong 3^2 is not equal to 6 so its \[\sqrt{84} - \sqrt{16}\] I think its \[2\sqrt{21} + 4 \] I am not quite sure but it can be simplified the two will not exist and 4 will be 2

OpenStudy (anonymous):

8

OpenStudy (lgbasallote):

?

OpenStudy (lgbasallote):

go on...keep shooting random guesses...im sure you'll hit the right one someday

OpenStudy (anonymous):

\[\sqrt{81}+3^2-6-\sqrt{16}\]=9+9-6-4=8

OpenStudy (anonymous):

if thats the question then the answer is 8 for sure.... :)

OpenStudy (anonymous):

Thank you so muchh :) hal_stirrup :D

OpenStudy (anonymous):

welcome @ngiee0118

OpenStudy (anonymous):

@Ingiee0118

OpenStudy (anonymous):

So your good at math ...? Can you help me with another question ? pleaseeeeeee .? :))

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

you can be my fan and so do i ?

OpenStudy (anonymous):

Which set of line segments with the given lengths form a right triangle? A ) 9, 27, 81 B ) 10, 24, 38 C )14, 16, 22 D) 15, 36, 39

OpenStudy (anonymous):

There i am a fan :)

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

So can you help me with that question ?

OpenStudy (anonymous):

its being some time since i did trig. i fine out.

OpenStudy (anonymous):

Huh ...?

OpenStudy (anonymous):

ok use pythagorean theorem, a^2 + b^2 = c^2

OpenStudy (anonymous):

Gur ... Okay thankyou (:

OpenStudy (anonymous):

the answer is D

OpenStudy (anonymous):

here is how its done. A ) 9, 27, 81 let 9=a , b=27 and c=81 the pythagorean theorem, a^2 + b^2 = c^2 so plug in the values and see if a^2 + b^2=c^2 .

OpenStudy (anonymous):

if this hold the you have the right answer. If not then try the B and c and D. THATS what i did and found out D IS the right answer

OpenStudy (anonymous):

Thankyouu(: So much !

OpenStudy (anonymous):

welcome. is that happy or sad face?

OpenStudy (anonymous):

HAPPY (:

OpenStudy (anonymous):

ok thats my happy face :)

OpenStudy (anonymous):

:D